# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

### v2

@@ IV-A Truncated SVD Adaptation
To gain deeper insights into the difference between LoRA and truncated SVD adaptation in FL, we introduce magnitude and directional discrepancies between the global model $\boldsymbol{\theta}_{\text{global}}$ and the local models $\boldsymbol{\theta}_{\text{local}}^{(i)}$ across selected clients. The magnitude discrepancy is defined as:
$\text{Mag}=\sum_{i=1}^{\mathcal{S}}\left\|\boldsymbol{\theta}_{\text{global}}-\boldsymbol{\theta}_{\text{local}}^{(i)}\right\|_{2},$ (5)
where $\mathcal{S}$ denotes the number of selected clients per round, and $\|\cdot\|_{2}$ represents the Frobenius norm. This metric captures the magnitude difference between the global and local models. The directional discrepancy is given by:
$\text{Dir}=\frac{1}{\mathcal{S}}\sum_{i=1}^{\mathcal{S}}\frac{\langle\boldsymbol{\theta}_{\text{global}},\boldsymbol{\theta}_{\text{local}}^{(i)}\rangle}{\|\boldsymbol{\theta}_{\text{global}}\|_{2}\,\|\boldsymbol{\theta}_{\text{local}}^{(i)}\|_{2}}.$ (6)
@@ IV-B1 Generation of Local Rank Masks
where $I(E_{n,i})$, $I(B_{n,ji})$, and $I(A_{n,ij})$ represent the importance scores of the corresponding parts. To avoid the massive overhead of sensitivity and gradient computation in AdaLoRA [32], the importance score $I(w_{i,j})$ in our design is directly represented in a simple yet effective way as $I(w_{i,j})=\left|w_{i,j}\right|$. As illustrated in Figure 6, each SVD module initially contains $r$ triplets, which can be inserted into six different components.
@@ IV-B2 Arbitration of Global Rank Masks
After collecting all candidates of the local rank masks from all selected clients, the server determines the global rank masks by arbitrating the fraction of clients reporting True at each position against a preset threshold $T_{h}$. This arbitration process is given by:
$M_{\mathrm{global}}^{(i)}=\begin{cases}True,&\text{if }\displaystyle\frac{1}{\lvert\mathcal{S}\rvert}\sum_{k\in\mathcal{S}}M_{k}^{(i)}>T_{h},\\ False,&\text{otherwise}.\end{cases}$ (9)
here, $\mathcal{S}$ denotes the set of participating clients. The arbitrated global rank masks are broadcast to all clients and used to optimize communication parameters between the server and clients.
@@ IV-B3 Communication with Arbitrated Masks
The global and local rank masks are transmitted alongside the parameters, yet their communication overhead remains negligible due to their boolean nature and significantly smaller size relative to the full parameters. During server-client communication, if a global rank mask in the truncated SVD modules is marked as False, the corresponding parameters are pruned. The complete weight matrix is reconstructed on the client or server using the global rank masks. This approach substantially reduces communication overhead in each round.
@@ V Experimental Setups
FedARA extends AdaLoRA [32], the SOTA dynamic parameter configuration method in centralized training, to an FL framework. We utilize Hugging Face’s Transformers library, which supports a wide range of pre-trained NLP models. FedARA uses a random client selection mechanism, which can be replaced with more sophisticated strategies [33, 34]. Communication is synchronized between clients and the cloud server using the weighted average method [7], ensuring compatibility with various optimization, and aggregation techniques [35, 36, 37, 38, 39, 40]. Additionally, FedARA supports advanced model quantization methods, such as QLoRA [41] and QA-LoRA [42], to further reduce computation and memory overhead while also potentially lowering communication costs.
Metrics The primary evaluation metric is the final accuracy and total communication overhead after a fixed number of FL rounds, ensuring consistent performance comparison. Secondary metrics include communication overhead per round, adaptive rank allocation results, total training time, memory usage, and energy consumption on edge devices. These metrics provide a comprehensive assessment of the system’s efficiency and effectiveness.
Hardware Following prior FL research [25], our experiments are conducted through emulation on a laptop with RTX 4070 GPU. The local training time is measured on three edge devices with hardware specifications similar to common resource-constrained devices, AGX Orin, Orin Nano, and Raspberry Pi 5. These training results were fed into our emulation framework to determine total training time in a homogeneous FL setting. In line with the existing works [25], we maintain a typical communication bandwidth of 1MB/s between server and clients. This setting accounts for a wide range of environments, from IoT to household WiFi and cellular networks.
TABLE I: Different dataset settings and sizes in sequence classification tasks of NLP.
Datasets Clients Num Labels Samples
20News [27] 100 20 18.8k
@@
Datasets and Division: FedARA is evaluated on four different NLP datasets, as detailed in Table I. (1) 20News [27]: content from 20 different newsgroups; (2) Semeval [43]: various NLP and semantic evaluation tasks; (3) AG News [44]: different types of news articles; (4) News Category [30]: a range of news categories. In the News Category dataset, we focus on the top 15 most frequent labels, covering 70% of the data. The datasets are split into training, validation, and testing sets in an 8:1:1 ratio. For IID settings, a Dirichlet distribution with $\alpha$ set to 1000. Non-IID settings follow FedAvg’s approach [7], partitioning each dataset by label, with each client receiving data for 1-2 labels, creating a pathological non-IID data distribution.
Model Our study employs two prominent NLP models: BERT [2] with 12 transformer blocks, and its streamlined variant DistilBERT [3] with 6 blocks, which retains near-original performance while offering a smaller model in size through knowledge distillation. Both models leverage the transformer architecture, which remains foundational for even the latest NLP models [45, 10]. Our research specifically targets federated adaptive allocation methods for transformer architectures. To ensure generality and feasibility on resource-constrained devices, we focus on BERT, DistilBERT, and sequence classification tasks in this paper, but our method is adaptable to emerging models and diverse tasks.
TABLE II: Comparison of FedPEFT methods across different models and datasets. In A (B), A represents the accuracy under non-IID data, while B denotes the relative accuracy degradation compared to IID data.
@@
FedLoRA-B-dr (r=12) 11.12 53.96 (↓10.99) 31.91 (↓48.46) 87.34 (↓1.55) 61.58 (↓10.14)
FedLoRA-B (r=6) 5.56 53.52 (↓11.24) 29.80 (↓50.29) 87.00 (↓1.70) 61.27 (↓10.36)
@@
FedLoRA-B-dr (r=12) 22.24 54.77 (↓9.15) 46.51 (↓34.16) 87.39 (↓1.68) 65.27 (↓6.91)
FedLoRA-B (r=6) 11.12 53.10 (↓11.33) 38.25 (↓42.23) 86.67 (↓1.45) 62.60 (↓9.81)
@@
Baselines We benchmark FedARA against several latest FedPEFT methods:
  • FedLoRA, which is a widely adopted FedPEFT method with extensive adaptations and studies [16, 17, 23, 19, 21, 22].
  • FedAdapter-h, which uses Adapter-h [11] as low-rank modules in both the attention and feed-forward network blocks of transformers.
  • FedAdapter-p, which inserts adapters only in the feed-forward network block [12]. AdaFL [25] is excluded from our experiments because it is a parameter-reduced variant of FedAdapter-p and lacks specific optimization for non-IID data.
  • FedLoRA-B, which is the SOTA method for mitigating non-IID effects [29, 24]. It exploits the distinct functions of LoRA matrices $B$ and $A$ by training only the more effective matrix $B$ while freezing $A$, thereby reducing interference from data heterogeneity and halving the number of trainable parameters.
  • FedLoRA-B-dr, the optimized method of FedLoRA-B, which trains only matrix B (corresponding to the matrix B in LoRA) but doubles the original rank to enhance its expressiveness while maintaining the total number of trainable parameters [28].
Hyperparameters Learning rates for different datasets and network models are determined through a grid search, ranging from 1e-5 to 1.5e-3. All methods use a batch size of 4, 1 epoch per FL round, a maximum sequence length of 64, and 100 rounds with 10 participating clients per round (10% and 1% selection rate). Adaptive rank allocation begins with 5 warm-up rounds, with rank decaying until the 50th round. The target rank is set to one-quarter of the original, with a global rank mask threshold of 0.5. Learning rates decrease linearly each round, using Adam as the optimizer. The scale hyperparameter alpha for LoRA is fixed at 16, following prior studies [15, 32].
@@ VI-A Performance under non-IID data
Table II illustrates the accuracy of various baselines across different datasets under non-IID, with FedARA’s initial rank $r$ set to 12 and 6, respectively. The best performance is highlighted in bold, while the second-best is underline. Each result is averaged over five different random seeds. The accuracy is presented in the format $A({\color[rgb]{0,0,1}B})$, where $A$ represents the accuracy under the non-IID setting, and $B$ denotes the decrease in accuracy relative to the IID setting, both expressed as percentages. A higher $A$ indicates better performance under the non-IID data, while a lower $B$ reflects less impact of non-IID on the method. We also list the total communication overhead measured in GigaBytes (GB).
@@
FedARA effectively mitigates the performance degradation caused by non-IID data. FedARA consistently achieves superior performance across four various datasets and two different models. Specifically, FedARA improves DistilBERT’s average accuracy on four datasets by 8.49% over FedLoRA and 11.58% over FedAdapter-h, while also surpassing FedLoRA-B-dr and FedLoRA-B by 6.95% and 7.75%, respectively. On the News Category dataset, FedARA enhances BERT’s performance by 9.78% compared to FedLoRA. Moreover, in comparison to FedAdapter-h, FedARA achieves its largest accuracy improvement of 37.59% on the Semeval dataset, which has many labels and a small sample size, while the smallest improvement of 1.47% is observed on the AG News dataset, which contains fewer labels and a large sample size. More importantly, FedARA significantly mitigates the effect of Non-IID data on the 20News and News Category datasets using BERT, reducing the performance gap from nearly 20% to 10% or less compared to FedLoRA.
The exceptional performance of FedARA is primarily attributed to its truncated SVD modules, which enhance similar feature representation across clients by enabling independent feature scaling and symmetric initialization. This method effectively mitigates the adverse effects of non-IID data.
Figure 7: Comparison of communication overhead per round on baselines (Model: BERT; Dataset: AG News).
@@
FedARA significantly improves communication efficiency. FedARA effectively reduces the communication overhead while maintaining superior performance compared to other baselines. As shown in Figure 7, despite starting with a similar initial communication cost (r=12) as baseline methods on the BERT model and AG News dataset, FedARA achieves a substantial reduction during federated fine-tuning, stabilizing at 131.42 MB, which is 71.17% lower than its initial cost during the warm-up round. Furthermore, as shown in Table II, when the initial rank of FedARA is twice that of FedLoRA, the total communication overhead of FedLoRA and FedLoRA-B-dr exceeds that of FedARA by nearly 20% across different models. When FedARA’s initial rank matches that of FedLoRA, the total communication overhead of FedLoRA and FedLoRA-B-dr is approximately 2.40$\times$ that of FedARA. Similarly, FedLoRA-B incurs nearly 20% more communication overhead than FedARA. Notably, FedARA achieves the lowest total communication cost while maintaining significantly better average performance compared to other baselines.
This improvement in communication efficiency stems from dynamic rank allocation, which continuously identifies and transmits the most critical parameters during each communication round while pruning less important ones. Figure 8 illustrates final rank allocation results for FedARA using the DistilBERT model and the 20News dataset. Initially set with a rank of 12 at each position, FedARA dynamically adjusts ranks to achieve a final average rank of 3, empirically balancing performance and communication efficiency. Darker blocks in Figure 8 indicate higher allocated ranks, primarily in deeper layers and key components such as query, key, f1, and f2, reinforcing their importance in retaining critical model information—consistent with our preliminary findings (Figure 2c).
@@
Figure 8: Results of final adaptive rank allocation (Model: DistilBERT; Dataset: 20News).

### v3

@@ IV-A Truncated SVD Adaptation
To further understand the difference in global weight updates between LoRA and truncated SVD adaptation in FL, we introduce magnitude and directional discrepancies between the global model $\boldsymbol{\theta}_{\text{global}}$ and the local models $\boldsymbol{\theta}_{\text{local}}^{(i)}$ across selected clients. The magnitude discrepancy is defined as:
$\text{Mag}=\sum_{i=1}^{K}\left\|\boldsymbol{\theta}_{\text{global}}-\boldsymbol{\theta}_{\text{local}}^{(i)}\right\|_{F},$ (9)
where $K$ denotes the number of selected clients per round, and $\|\cdot\|_{F}$ represents the Frobenius norm. This metric captures the magnitude difference between the global and local models. The directional discrepancy is given by:
$\text{Dir}=\frac{1}{{K}}\sum_{i=1}^{{K}}\frac{\langle\boldsymbol{\theta}_{\text{global}},\boldsymbol{\theta}_{\text{local}}^{(i)}\rangle}{\|\boldsymbol{\theta}_{\text{global}}\|_{F}\,\|\boldsymbol{\theta}_{\text{local}}^{(i)}\|_{F}}.$ (10)
@@ IV-B1 Generation of Local Rank Masks
where $I(E_{n,i})$, $I(B_{n,ji})$, and $I(A_{n,ij})$ denote the importance scores of the corresponding parts. As shown in Table I, we compare four importance score evaluation strategies: Mag denotes the weight magnitude, $I(w_{i,j})=|w_{i,j}|$; Grad represents the gradient, $|\text{grad}(w_{i,j})|$; Mixed combines both weight and gradient information via their product, $|\text{grad}(w_{i,j})\cdot w_{i,j}|$; and Sensitivity corresponds to the sensitivity-based method proposed in AdaLoRA [28]. Among the evaluated strategies, the magnitude-based method consistently achieves the best performance across different datasets and models, outperforming both the mixed strategy and gradient-based approaches. While the sensitivity-based method also improves accuracy, it introduces approximately 30% additional computational overhead. To avoid the high cost of sensitivity and gradient-based estimation, we adopt the magnitude-based score as a simple yet effective alternative.
TABLE I: Performance and local computation overhead under different importance score evaluation methods (non-IID alpha=0.1). Results are reported in percentage (%).
Methods DistilBERT BERT Comp.
20News Semeval 20News Semeval
Mag 63.89 80.24 64.32 76.89 1$\times$
Grad 63.31 79.59 63.68 76.05 0.97$\times$
Mixed 63.79 79.87 63.94 76.42 1.11$\times$
Sensitivity 63.89 80.06 64.21 76.70 1.30$\times$
@@
TABLE II: Performance and communication overhead under different arbitration strategies (non-IID alpha=0.1). Results are reported in percentage(%).
Methods 20News Comm.
DistilBERT BERT
FedARA 62.94 63.52 8.80 GB
FedARA-global 62.78 63.36 8.82 GB
@@ IV-B2 Arbitration of Global Rank Masks
After collecting all candidates of local rank masks from all selected clients, the server determines global rank masks by arbitrating the fraction of clients reporting True at each position against a preset threshold $T_{h}$. This arbitration process is given by:
$M_{\mathrm{global}}^{(i)}=\begin{cases}True,&\text{if }\displaystyle\frac{1}{\lvert{K}\rvert}\sum_{k\in{K}}M_{k}^{(i)}>T_{h},\\ False,&\text{otherwise}.\end{cases}$ (13)
here, ${K}$ denotes the set of participating clients. The arbitrated global rank masks are broadcast to all clients and used to optimize communication parameters between the server and clients. As shown in Table II, we compare two arbitration strategies: FedARA collects local rank masks from clients and performs arbitration on the server; and FedARA-global generates global rank masks directly from the aggregated model. Although the communication overhead is similar for both strategies, FedARA achieves better performance on the 20News dataset [25], highlighting the effectiveness of leveraging local rank information for centralized arbitration.
@@ IV-B3 Communication with Arbitrated Masks
The global and local rank masks are transmitted alongside the parameters, yet their communication overhead remains negligible due to their boolean nature and significantly smaller size relative to the full parameters. During server-client communication, if a global rank mask in the truncated SVD modules is marked as False, the corresponding parameters are pruned. The complete weight matrix is reconstructed on the client or server using global rank masks. This approach substantially reduces communication overhead in each round.
@@ V Experimental Setups
FedARA extends AdaLoRA [28], originally designed for dynamic parameter configuration in centralized training, to the federated learning setting. Built on Hugging Face’s Transformers library, it supports a broad range of pre-trained NLP models. FedARA adopts a random client selection mechanism, which can be substituted with more advanced strategies [29, 30]. Communication between clients and the server is synchronized via weighted averaging [7], maintaining compatibility with diverse optimization and aggregation methods [31, 32, 33, 34, 35, 36]. Moreover, FedARA incorporates model quantization techniques such as QLoRA [37] and QA-LoRA [38] to reduce computational and memory overhead, potentially lowering communication costs.
Metrics. The primary evaluation metrics are final accuracy for classification tasks, ROUGE scores [39] for summarization tasks, and total communication overhead after a fixed number of FL rounds, ensuring fair performance comparison. Secondary metrics include per-round communication cost, adaptive rank allocation, total training time, memory footprint, and energy consumption on edge devices. These metrics offer a comprehensive evaluation of system efficiency and effectiveness.
Hardware. Following prior FL research [23], all experiments are conducted via emulation on a laptop equipped with an RTX 4070 GPU. Local training time is measured on three edge devices—AGX Orin, Orin Nano, and Raspberry Pi 5—chosen for their representativeness of common resource-constrained hardware. These results are integrated into our emulation framework to estimate total training time in a homogeneous FL setting. In line with previous studies [23], we assume a communication bandwidth of 1MB/s between the server and clients, reflecting realistic conditions across IoT, home WiFi, and cellular networks.
TABLE III: Different dataset settings and sizes in sequence classification and sumarization tasks of NLP.
Tasks Datasets Client Num. Labels Samples
Classification 20News [25] 100 20 18.8k
@@
Summarization CNN/DailyMail [42] 1000 - 287k
Datasets and Division. FedARA is evaluated on sequence classification and summarization tasks using five diverse NLP datasets, as summarized in Table III. Classification tasks include: (1) 20News [25], comprising documents from 20 different newsgroups; (2) Semeval [40], covering various NLP and semantic evaluation tasks; (3) AG News [41], containing news articles from four categories; (4) News Category [27], a dataset of diverse news headlines, where we focus on the top 15 most frequent labels, covering approximately 70% of the data. Summarization task: (5) CNN/DailyMail [42], a widely used dataset for abstractive summarization. Each classification dataset is split into training, validation, and test sets in an 8:1:1 ratio. We fine-tune a DistilBERT model on the News Category dataset (top-15 labels) and use the resulting model to generate pseudo-labels for the summarization dataset. To simulate non-IID conditions in a federated setup, we adopt Dirichlet-based partitioning with $\alpha\in\{1,0.1,0.01\}$. Following FedAvg [7], we also consider a pathological non-IID setting where each client holds samples from only 1–2 labels. For IID settings, we set $\alpha=1000$ to approximate uniform label distribution across clients.
Model. We adopt three widely used transformer-based models for NLP tasks. For classification, we use BERT-base [3], which consists of 12 transformer layers, and its compressed variant DistilBERT-base [2], which retains most of BERT’s performance with only 6 layers via knowledge distillation. For summarization, we employ BART-base [4], an encoder–decoder model designed for sequence-to-sequence generation. All selected models are based on the transformer architecture, which remains foundational for even the latest NLP models [10]. Our work focuses on federated adaptive rank allocation for transformer models. While this study evaluates BERT, DistilBERT, and BART across classification and summarization tasks to ensure generality and feasibility on resource-constrained devices, our method is readily applicable to newer architectures and broader NLP tasks.
TABLE IV: Comparison of FedPEFT methods across different models and datasets. In A (B), A represents the accuracy under non-IID data (Pathological distribution), while B denotes the relative accuracy degradation compared to IID data.
@@
SLoRA (r=6) 11.12 48.83 (↓15.33) 31.74 (↓47.01) 85.68 (↓2.17) 55.57 (↓14.91)
FeDeRA (r=6) 11.12 52.28 (↓11.43) 28.33 (↓50.95) 86.22 (↓1.94) 56.45 (↓14.64)
FFA-LoRA-dr (r=12) 11.12 53.96 (↓10.99) 31.91 (↓48.46) 87.34 (↓1.55) 61.58 (↓10.14)
FFA-LoRA (r=6) 5.56 53.52 (↓11.24) 29.80 (↓50.29) 87.00 (↓1.70) 61.27 (↓10.36)
@@
SLoRA (r=6) 22.24 45.60 (↓16.73) 28.42 (↓47.93) 84.36 (↓3.37) 54.20 (↓15.76)
FeDeRA (r=6) 22.24 50.58 (↓13.95) 31.69 (↓49.02) 86.97 (↓2.42) 54.82 (↓17.72)
FFA-LoRA-dr (r=12) 22.24 54.77 (↓9.15) 46.51 (↓34.16) 87.39 (↓1.68) 65.27 (↓6.91)
FFA-LoRA (r=6) 11.12 53.10 (↓11.33) 38.25 (↓42.23) 86.67 (↓1.45) 62.60 (↓9.81)
@@
Baselines. We compare FedARA against several recent FedPEFT methods:
  • FedLoRA: a widely adopted baseline with extensive variants and empirical studies [16, 17, 21, 18].
  • FedAdapter-h: incorporates Adapter-h [11] modules into the attention and feed-forward layers of transformers.
  • FedAdapter-p: inserts adapter modules only in the feed-forward layers [12]. We exclude AdaFL [23] as it is a IIIparameter-reduced version of FedAdapter-p without targeted non-IID optimization.
  • SLoRA: employs a two-stage training scheme [19], where the first stage applies sparse fine-tuning (SFT) to improve initialization for LoRA in the second stage. We allocate 10% of the total FL rounds to the SFT stage.
  • FeDeRA: initializes LoRA using singular value decomposition of pre-trained weights [20].
  • FFA-LoRA: updates only the $B$ matrix in LoRA while freezing $A$ [26, 22, 43]. To mitigate reduced expressiveness, we also evaluate FFA-LoRA-dr, which employs orthogonal initialization for $A$ and doubles the rank.
Hyperparameters. Learning rates are selected via grid search in the range of 1e-5 to 5e-3, depending on the dataset and model. All methods use a batch size of 4, one epoch per FL round, a maximum sequence length of 64, and are fine-tuned for 100 rounds with 10 clients per round (10% or 1% selection). For summarization, we set the source length to 512, the target length to 64, and the beam size to 1 to fit edge device constraints. Adaptive rank allocation starts after 5 warm-up rounds and decays until round 50, targeting one-quarter of the original rank with a global mask threshold of 0.5. Learning rates decay linearly across rounds, using the Adam optimizer. The LoRA scaling factor $\alpha$ is fixed at 16, following prior work [15, 28].
@@ VI-A Performance under non-IID data
We conduct experiments on classification tasks under pathological non-IID and IID settings, using two different models across four datasets. The results are summarized in Table IV, where the best performance is highlighted in bold and the second-best is underlined. Each result is averaged over five runs with different random seeds.
@@
FedARA effectively mitigates performance degradation under non-IID settings. FedARA consistently outperforms existing methods across four datasets and two models. Using DistilBERT, FedARA achieves an average accuracy improvement of 8.49% over FedLoRA and 11.58% over FedAdapter-h, while also surpassing FFA-LoRA-dr and FFA-LoRA by 6.95% and 7.75%, respectively. The largest gain 37.59% is observed on the Semeval dataset, which has more labels and fewer samples, whereas the smallest gain 1.47% occurs on AG News, with fewer labels and more samples. Moreover, FedARA significantly outperforms SLoRA and FeDeRA, which rely on modified LoRA initialization strategies.
We further conduct experiments under varying degrees of non-IID data distributions by adjusting the Dirichlet parameter $\alpha$, where smaller $\alpha$ indicates higher non-IIDness. The results are illustrated in Figure 7. As $\alpha$ decreases, the performance of FedPEFT shows a clear downward trend, which is consistent with the observations reported in [19, 20]. Moreover, FedARA consistently outperforms other FedPEFT methods across different $\alpha$ values, with the performance gap becoming particularly significant under highly skewed non-IID settings.
Within our experimental framework, SLoRA and FeDeRA underperform FedARA on certain non-IID tasks and datasets. SLoRA requires careful hyperparameter tuning during its initial sparse fine-tuning stage, especially under varying non-IID data distributions and tasks. Its advantage tends to diminish as LoRA’s rapid adaptation under the Adam optimizer. The effectiveness of FeDeRA relies on alignment between the pre-trained model and the downstream task, which becomes increasingly difficult to maintain under non-IID conditions, potentially leading to performance degradation when such alignment is weak. In contrast, FedARA leverages a scaling vector and symmetric initialization to reduce client drift and promote consistent feature representations, thereby mitigating performance degradation.
Figure 7: Accuracy comparison under different data distributions, with parameter settings identical to those in Table IV.
@@
FedARA significantly improves communication efficiency. FedARA reduces the communication overhead while maintaining superior performance. As shown in Figure 8, FedARA begins with a similar initial communication cost (r=12) as baseline methods on the BERT model and AG News dataset, but quickly drops it to 131.42 MB—71.17% lower than its initial warm-up cost. SLoRA and FeDeRA incur the same communication cost as FedLoRA. Table IV further shows that even when FedARA’s initial rank is twice that of FedLoRA, its total communication overhead remains nearly 20% lower than that of FedLoRA and FFA-LoRA-dr. When initial ranks are equal, FedARA reduces communication overhead by approximately 2.40$\times$. FFA-LoRA also incurs about 20% more overhead than FedARA. Overall, FedARA achieves the lowest communication cost across settings while consistently outperforming all baselines.
Figure 8: Comparison of communication overhead per round across baselines.
@@
Figure 9: Results of final adaptive rank allocation using the DistilBERT model.
Figure 9 shows the final rank distribution of FedARA on the DistilBERT model and 20News dataset. Starting from a uniform rank of 12, FedARA adaptively reduces the average rank to 3, effectively balancing performance and communication cost. Darker blocks in Figure 8 represent higher ranks, mainly in deeper layers and key components like query, key, f1, and f2, highlighting their role in preserving important model information—consistent with our earlier findings (Figure 2c).
To evaluate the generalization and scalability of FedARA, we conduct summarization experiments using the BART model on the CNN/DailyMail dataset. FedARA demonstrates strong performance while achieving the best communication efficiency among all compared methods. As shown in Table V, FedARA improves ROUGE scores over FFA-LoRA, FFA-LoRA-dr, and FedLoRA. FedARA delivers performance comparable to SLoRA and FeDeRA while reducing communication overhead by about 60%. These results confirm its effectiveness in generative NLP tasks with encoder-decoder architectures, in addition to classification.
TABLE V: Results on text generation tasks fine-tuned by the BART [4] model for different baselines. non-IID alpha is set to 0.1.
Datasets Methods Comm. (GB) ROUGE 1/2/L (%)
CNN /DailyMail FedLoRA (r=12) 27.19 40.79/18.33/28.70
SLoRA (r=12) 27.19 40.90/18.51/28.98
FeDeRA (r=12) 27.19 40.90/18.49/29.00
FFA-LoRA-dr (r=24) 27.19 40.71/18.28/28.66
FFA-LoRA (r=12) 13.60 40.58/18.14/28.55
FedARA (init r=12) 9.77 40.93/18.49/29.00